Given the production function:

 Y = 10 + 5X1 – 0.2 X12

a)  Define the functions for average product and marginal product.

b)  Compute the maximum level of production of Y and the corresponding    level of input X1.

c)        Given the price of Y = R3.75/unit and that of X1 = R1.50/unit, algebraically compute the maximum profit level of input and output.      


1
Expert's answer
2021-06-03T09:12:09-0400

a)


"AY=\frac{10}{X_1}+5-0.2X_1"

"MY=5-0.4X_1"

b)


"5-0.4X_1=0"

"X_1=12.5"

"Y=41.25"

c)


"\pi=41.25\times3.75-12.5\times1.5=135.94"